gtt ekn Manual

Edition : 2016.03.28

Y.Goto, Y.Tachibana, N.Takayama

1 Function Manual

1.1 Outline

1.2 Notation

1.3 Functions

1.3.1 gtt_ekn.nc
gtt_ekn.nc(beta,p)

:: It returns the value of the normalizing contant and its derivatives of the conditional hypergeometric distribution of two way contingecy tables with fixed marginals

return [Z,[[d 11 Z, d 12 Z, ...], ..., [d m1 Z, d m2 Z, ...., d mn Z]]]

beta List of row sums and column sums. All entries must be positive integer.

p The probabilities of the cells of the table.

• Details have not been written. See Japanse documents.

2x3 contingency table with the row sum [4,5] and the column sum [2,4,3].

     [2237] gtt_ekn.nc([[4,5],[2,4,3]],[[1,1/2,1/3],[1,1,1]]);
     [4483/124416,[ 353/7776 1961/15552 185/1728 ]
     [ 553/20736 1261/15552 1001/13824 ]]

English translation for other functions has not yet been written.

Index 2

Index

(Index is nonexistent) (Index is nonexistent)

Short Contents

1 Function Manual. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

Table of Contents

1 Function Manual. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Outline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Notation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.3.1 gtt_ekn.nc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2